OpenStudy (anonymous):

The equation of a circle is shown below. (x – 3)2 + (y – 2)2 = 9 Which statement is true about the center of the circle? It is located on the y-axis. It is located in the third quadrant. It is located in the first quadrant. It is located on the x-axis.

jimthompson5910 (jim_thompson5910):

Do you have a graphing calculator?

jimthompson5910 (jim_thompson5910):

Or graph paper?

OpenStudy (anonymous):

i have geogebra on my computer!

jimthompson5910 (jim_thompson5910):

very nice, so do I, and I love it lol

jimthompson5910 (jim_thompson5910):

simply type (x-3)^2 + (y-2)^2 = 9 into the input bar and it will graph the circle

OpenStudy (anonymous):

lol me to!!! ...it doesnt do a circle tho it does a line

jimthompson5910 (jim_thompson5910):

remember to type ^

jimthompson5910 (jim_thompson5910):

It's NOT (x – 3)2 + (y – 2)2 = 9 it's really (x-3)^2 + (y-2)^2 = 9

OpenStudy (anonymous):

oo ok hold on ill redo it! lol

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

ok now i have a circle! lol

jimthompson5910 (jim_thompson5910):

yay

jimthompson5910 (jim_thompson5910):

so where is the center?

OpenStudy (anonymous):

is this right??? ....It is located in the first quadrant.

jimthompson5910 (jim_thompson5910):

optionally, you can type "center(c)" without quotes and it will plot out the center point

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

wait no ...ummm this one right?? ....It is located on the x-axis.

jimthompson5910 (jim_thompson5910):

no, did you type "center(c)"?

jimthompson5910 (jim_thompson5910):

into the input bar

jimthompson5910 (jim_thompson5910):

oh and don't enter the quotes

OpenStudy (anonymous):

i didnt know it could do that! lol ...center is (3,2)

jimthompson5910 (jim_thompson5910):

perfect

jimthompson5910 (jim_thompson5910):

btw, if you can't use a graphing calculator or geogebra, then you can look at the equation (x-3)^2 + (y-2)^2 = 9 and notice that it's in (x-h)^2+(y-k)^2=r^2 form Since h = 3 and k = 2, the center is (3,2)

OpenStudy (anonymous):

ok...so whats the answer then?

jimthompson5910 (jim_thompson5910):

where is (3,2)? is it on the x-axis or in the first quadrant?

OpenStudy (anonymous):

ummm first quadrant??

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

so the center is in the first quadrant

jimthompson5910 (jim_thompson5910):

your first answer was right on, so it's always good to go with your instinct

OpenStudy (anonymous):

ok yay! lol ...thanks!!!!

jimthompson5910 (jim_thompson5910):

yw